# Feature Request Analysis and Processing: Tasks, Techniques, and Trends

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Requirements analysis
@@ 2.2. Related Work
Although several secondary studies have investigated concepts related to user feedback and software features, most have focused on adjacent topics or partial perspectives rather than offering a systematic synthesis of feature request research across the software development lifecycle. Existing work has examined areas such as change request repositories (Cavalcanti et al., 2014), requirement or information extraction from app reviews (Tavakoli et al., 2018; Cai et al., 2025), feature modeling (Bischoff et al., 2019), feature extraction in software product lines (Bakar et al., 2015), feature localization (Dit et al., 2013), and high-level mapping studies on crowdsourced feedback (Wang et al., 2019; Dąbrowski et al., 2022). However, these studies do not provide a unified view of how feature requests are managed throughout the software lifecycle.
@@ 3.2.2. Inclusion/Exclusion Criteria
To identify the most relevant studies on feature request analysis and processing, we drew upon prior work (Giray et al., 2023; Zakari et al., 2020; Croft et al., 2022; Niu et al., 2025b; Hu et al., 2025) and established a set of inclusion (ICs) and exclusion criteria (ECs), summarized in Table 1. These criteria were applied to the titles, abstracts, and keywords of candidate papers to ensure relevance and quality.
@@ 3.3. Data Extraction and Synthesis
    • Data preprocessing techniques and feature extraction methods are described.
@@
The goal of this survey is to aggregate evidence from primary studies and provide an overview of the state of the art on feature request analysis and processing. Following the systematic review guidelines (Keele et al., 2007), we synthesized the extracted data using a combination of descriptive (quantitative) synthesis and qualitative thematic synthesis, aligned with our four research questions. For RQ1 (Demographics), we performed descriptive synthesis on bibliographic and contextual attributes (e.g., publication year, venue type, and study type). We report the evolution of the literature over time and the distribution across venues using frequency counts. For RQ2 (Analysis and Processing Practices), we synthesized the extracted methodological information by grouping studies according to the main feature request tasks they target (e.g., classification, prioritization, summarization, management) and the approaches they employ. We summarize the prevalence of different practices (including preprocessing, feature extraction/representation, and evaluation setups) using structured comparison tables, and we narratively discuss commonalities and differences across studies. For RQ3 (Public Datasets and Tools), we first summarized availability-related information (e.g., whether datasets/tools are publicly accessible, their sources) in a structured catalog. We then conducted a qualitative categorization of the datasets and tools by their intended purposes and research themes, and reported results for each category to facilitate reuse. For RQ4 (Challenges and Opportunities), we conducted an inductive thematic synthesis. During data extraction, any text explicitly describing a limitation, difficulty, research gap, or future work was copied verbatim into the extraction form. Two reviewers independently performed open coding on these excerpts to identify recurring concepts and iteratively refined a shared codebook through discussion. Conceptually similar codes were merged into higher-level themes, which were manually reviewed to ensure consistent granularity and clear boundaries. We report the resulting themes together with representative evidence from primary studies and, where appropriate, the number of studies mentioning each theme. When divergent or conflicting observations emerged, we explicitly described them rather than forcing a single conclusion.
@@ 4. Demographics
A total of 140 research papers were included in the final study set, covering publications up to August 2025. As shown in Figure 4, the number of primary studies has increased significantly over time, showing a clear upward trend from 2011 to 2017, with the peak occurring in 2017 and 2019 (16 studies). Activity remained relatively high through 2020, followed by a decline likely due to shifting research priorities. A notable resurgence occurred in 2024 (11 studies), coinciding with the rise of LLMs, which introduced new research opportunities. Overall, the data reveals cyclical research activity with notable peaks and dips over the years. Figure 5 presents the distribution of the selected papers across journals and conferences, highlighting the most frequently appearing publication venues. A total of 88 papers were published in conferences, 46 in journals, and 6 as preprints. Most studies were published in top-tier software engineering outlets, particularly in the field of requirements engineering. High-frequency venues include International Conference on Requirements Engineering (RE Conf., 15 papers), International Conference on Automated Software Engineering (ASE Conf., 8 papers), International Working Conference on Requirement Engineering: Foundation for Software Quality (REFSQ, 7papers), International Conference on Software Engineering (ICSE, 7 papers), Empirical Software Engineering (ESE, 7 papers), Journal of Systems and Software (JSS, 6 papers), Transactions on Software Engineering (TSE, 4 papers), International Conference on Software Maintenance and Evolution (ICSME, 4 papers), Requirements Engineering Journal (RE Jour., 4 papers), and International Conference on Evaluation and Assessment in Software Engineering (EASE, 4 papers). These results confirm that the topic has received strong attention from the software engineering community.
@@ 5.2. Quality Review
Quality review is a critical activity in requirements engineering, aimed at assessing the quality, correctness, and feasibility of incoming requirements. As noted by Wiegers (Wiegers, 2003), a comprehensive review should examine aspects such as the organization, completeness, correctness, quality attributes, and traceability of the requirements specification. Functioning as a “filter” within the requirements analysis process, feature request review helps eliminate low-quality requests—such as those that are ambiguous, redundant, or irrelevant—before they reach the development phase, thereby reducing errors and defects downstream (Wiegers, 2003).
@@ 5.4.1. Lexical Analysis-Based Approaches
Regarding priority ranking and competitive analysis, Licorish et al. (Licorish et al., 2017) focused on noun extraction from reviews to predict the processing priority of feature requests, while Dalpiaz et al. (Dalpiaz and Parente, 2019) utilized POS tagging and various POS combinations—such as ¡noun + adjective¿ and ¡adjective + verb¿—to mine features from competitor app reviews. Shah et al. (Shah et al., 2018) proposed annotation guidelines emphasizing the labeling of only consecutive word phrases containing nouns, limited to a maximum of three words, while excluding references to the application itself.
@@ 5.7. Classification
✎ Classification Categories. Regarding classification categories, feature requests are typically organized according to requirement levels described in the requests, such as security, usability, ease of use, among others. Requirements engineering commonly distinguishes between functional requirements—detailing software functionalities—and non-functional requirements, which specify quality attributes essential for fulfilling user and business needs. Standards like ISO 25010 categorize software quality into functional and non-functional characteristics, with the latter encompassing functional suitability, performance efficiency, compatibility, usability, reliability, security, maintainability, and portability (ISO/IEC, 2001). Various studies have classified feature requests accordingly. Li et al. (Li et al., 2018) categorized requests into security, reliability, performance, lifecycle, usability, ease of use, and system interface. Niu et al. (Niu et al., 2021) adopted the same dataset as Li et al. Scalabrino et al. (Scalabrino et al., 2019) categorized user comments into defect reports, functional requests, performance requests, security requests, energy consumption requests, availability requests, and other categories. Tao et al. (Tao et al., 2020) classified security-related comments into system, finance, spam, privacy, and others. Wang et al. (Wang et al., 2018) grouped requirements from user reviews into usability, reliability, portability, and performance. More recently, Bai et al. (Bai et al., 2024) used structural topic modeling to identify eight key user concerns from video conferencing app reviews, including data privacy, audio/video quality, customer service, security issues, meeting management, feature requests, and platform compatibility.
In conclusion, feature request classification is a crucial step in requirements analysis that facilitates effective organization, understanding, and management of diverse user requests. The field has evolved from early unsupervised clustering techniques to more advanced supervised learning and explainable AI approaches, leveraging both textual and auxiliary features for improved accuracy. Categorization schemes generally align with established requirements engineering frameworks, distinguishing functional and non-functional aspects and capturing detailed quality attributes. Continued advancements in classification methodologies and feature representation promise to further enhance automated analysis and prioritization of feature requests in software development.
@@ 5.9.1. Horizontal Traceability of Feature Requests
Horizontal traceability involves creating connections among software artifacts that exist at the same abstraction level or within a single development phase, particularly within requirements engineering (Heck and Zaidman, 2014). For feature requests, establishing horizontal traceability aids developers and stakeholders in gaining clear insights into project progress, efficiently managing the integration of new features, detecting requirement conflicts at an early stage, comprehending the rationale behind software requirements, handling interdependencies logically, and avoiding redundant or contradictory tasks, thereby reducing risks of software project failures.
@@ 5.11. Code Localization
One of the earliest and most influential studies in this field is by Marcus et al. (Marcus et al., 2004), who introduced an IR-based approach using LSI for concept and feature localization. The method represents source code entities as textual documents and computes their semantic similarity with feature request descriptions in a latent space, enabling the identification of related code even without exact keyword matches. This seminal work laid the foundation for later IR-based and hybrid localization techniques.
@@ 6.2. Publicly Available Datasets
A significant impediment in many areas of requirements engineering research is the scarcity of publicly accessible datasets, often due to the proprietary nature of software requirements (Yu et al., 2024). However, the domain of feature requests presents a unique advantage: a substantial volume of data originates from open platforms, such as public repositories and app stores. This accessibility, often leveraged through web scraping techniques, has partially alleviated the data scarcity challenge that plagues other RE sub-fields. Consequently, various datasets have been curated and made publicly available by researchers. Table LABEL:tab:dataset_long catalogs these resources, detailing their origins and the specific research tasks they are typically used for (e.g., training classification models, evaluating duplicate detection).
@@
(Iacob and Harrison, 2013) http://goo.gl/cL00vd Identify feature requests from user review. 2013
(Maalej and Nabil, 2015) https://mast.informatik.uni-hamburg.de/app-review-analysis/ 2015
(Maalej et al., 2016) https://mast.informatik.uni-hamburg.de/app-review-analysis/ 2016
(Wang et al., 2023) https://github.com/KeyL99/VoteBot/ 2023
@@
(Licorish et al., 2015) https://goo.gl/4qLVcf Prioritization. 2015
(Keertipati et al., 2016) https://play.google.com/store/apps/details?id=com.google.android.maps.mytracks 2016
(Scalabrino et al., 2019) https://dibt.unimol.it/reports/clap/ 2017
@@
(Fitzgerald et al., 2012) http://sre-research.cs.ucl.ac.uk/Intueri/ Failure prediction. 2012
(Palomba et al., 2017) https://sites.google.com/site/changeadvisormobile/ Feature localization. 2017
@@
(Hu et al., 2022) https://github.com/lcynju/review2code Feature localization. 2022
@@ 7.1. Summarization
Before delving into the core findings, the study established essential context by defining software features, feature requests, and related concepts, underscoring their value and distinguishing them from traditional, formally elicited software requirements. We also positioned this review in relation to existing surveys, clarifying its unique contribution in providing an end-to-end requirements engineering perspective on the feature request lifecycle (Section 2).
The main body of our findings (presented from Section 4 onwards) begins by outlining the demographics of the research landscape. This revealed a consistent growth in publications dedicated to feature requests since 2010, with significant activity concentrated in reputable software engineering and requirements engineering venues, confirming the topic’s increasing importance within the community. We systematically cataloged the diverse origins of feature requests, confirming that while dedicated platforms like issue trackers are used, a vast amount originates from less structured channels such as app reviews, social media, and developer communications, with app reviews and issue reports being the most frequently studied sources (Section 5.1).
Addressing RQ3, the core analysis examined how these requests are analyzed and processed within a requirements engineering framework. We reviewed various automated methods designed to identify feature request descriptions from often noisy and informal user feedback, spanning unsupervised, supervised, and semi-supervised paradigms. Then we discussed techniques for software feature extraction, often relying on lexical approaches like POS tagging. It also included methods for structural analysis aimed at understanding content, such as user intent recognition, classification into diverse functional or non-functional categories, and automated summarization. Furthermore, the critical task of prioritization was examined, contrasting traditional RE approaches with methods adapted specifically for the high volume and nature of user feedback. Moreover, we examined approaches for quality review, primarily focusing on quality assessment related to duplication, redundancy, and ambiguity. We also explored traceability methods, encompassing both horizontal links between related requests and vertical links to artifacts such as source code and developers. Finally, various facets of change management were reviewed, including predicting request approval, recommending suitable developers, localizing relevant code sections, and suggesting APIs or refactorings.
@@ 7.2. Challenges and Opportunities
Comparative Analysis of Feature Requests and Bug Reports. Firstly, this review has focused exclusively on feature requests. However, user feedback often intertwines requests for new functionality with reports of existing problems. A crucial future direction is to integrate the study of bug reports alongside feature requests. Adopting a requirements engineering lens, comparative analyses are needed to explore the similarities and differences in effectively analyzing, prioritizing, and managing these two distinct yet interconnected types of feedback. Understanding where techniques converge and where they must diverge could lead to more holistic and efficient feedback management frameworks and tooling.
@@
Specification and Validation of Feature Requests. Furthermore, a significant and largely unaddressed gap identified in our study pertains to the specification and validation of feature requests. Currently, research heavily emphasizes identification and analysis, but largely overlooks the critical subsequent steps needed to bridge the gap between raw user input and actionable, verifiable requirements. Future research must urgently tackle the challenge of transforming collections of informal, often ambiguous, and fragmented user requests into more structured, consistent, and usable requirement representations. This could involve exploring techniques for semi-automatically generating preliminary user stories, use case scenarios, behavioral models, or even identifying constraints and non-functional aspects implicitly mentioned in feedback. Equally vital, and even less explored, is the validation of these derived specifications. Methodologies and practical techniques are needed to ensure that the interpretations and specifications derived from user feedback accurately reflect the original user needs and intent before substantial development effort is invested. Addressing this specification and validation gap is fundamental to truly leveraging feature requests for building software that meets genuine user expectations. We plan to conduct in-depth investigations into these specific areas in our future work.
@@
Leveraging LLMs for Feature Request Intelligence. With the rapid advancement of LLMs, their application in software engineering tasks has gained increasing traction. However, our review finds that feature request research has yet to leverage the potential of LLMs fully. These models offer promising capabilities for a wide range of tasks, including classifying user intent, mapping feature requests to relevant source code locations, summarizing feedback, and recommending suitable developers or implementation strategies. Future work should explore how LLMs can be fine-tuned or prompted for domain-specific use in feature request processing, and investigate their effectiveness compared to traditional machine learning techniques. Moreover, new evaluation criteria and human-in-the-loop workflows may be necessary to ensure reliability, interpretability, and alignment with user expectations.
@@
Lack of Benchmarks for LLM-Driven Feature Request Tasks. Despite growing interest in applying LLMs to software maintenance and requirements tasks, there is a notable absence of standardized benchmarks specifically targeting LLM-based feature request processing. This limits reproducibility, fair comparison across methods, and progress tracking over time. To address this, the research community should work toward building and releasing well-curated, multi-task benchmark datasets that reflect real-world feature request scenarios. These benchmarks should include annotations for classification, prioritization, code mapping, and developer assignment, and be accompanied by baseline LLM implementations to support further development and evaluation.
@@
Multilingual and Cross-Cultural Feature Request Processing. Current research on feature requests is heavily skewed toward English-language feedback, predominantly from Western platforms and user bases. However, in today’s global software market, products often receive feedback in multiple languages and cultural contexts. This introduces new challenges in understanding linguistic nuances, implicit user expectations, and regional preferences. Future work should explore multilingual models and culturally adaptive techniques to support accurate identification and interpretation of feature requests across diverse user groups. This direction also calls for the development of multilingual corpora and evaluation tools.
@@
Bridging User and Developer Perspectives. Feature requests often originate from non-technical users but are consumed and implemented by developers, creating a semantic and expectation gap. Bridging this gap remains an underexplored area. Future work could investigate how to translate user-facing language into developer-understandable representations, integrate user sentiment and urgency into development planning, and develop tools that facilitate two-way communication between stakeholders. Enhancing alignment between user needs and developer actions could significantly improve the effectiveness and user satisfaction of feature implementation.
Autonomous Agents for End-to-End Feature Request Processing. Another emerging and promising direction for future research is the development of autonomous agents capable of handling feature requests throughout the entire lifecycle—from initial user interaction and intent interpretation to analysis, code localization, and implementation. With advancements in multi-agent collaboration frameworks and AI planning, it is increasingly feasible to envision intelligent agents that can process user feedback, generate actionable development tasks, and even coordinate with other agents for implementation and testing. Such agent-based systems could significantly reduce the human effort required for managing feature requests, especially in large-scale and user-driven software systems. Exploring the design, capabilities, and limitations of such autonomous feature request agents offers a valuable avenue for advancing intelligent software engineering.

### v3

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Requirements analysiscopyright: ccjournal: TOSEMjournalyear: 2026journalvolume: 1journalnumber: 1publicationmonth: 1doi: 10.1145/3799899
[…]
by
@@ 2.2. Related Work
Although several secondary studies have investigated concepts related to user feedback and software features, most have focused on adjacent topics or partial perspectives rather than offering a systematic synthesis of feature request research across the software development lifecycle. Existing work has examined areas such as change request repositories (Cavalcanti et al., 2014), requirement or information extraction from app reviews (Tavakoli et al., 2018; Cai et al., 2025), feature modeling (Bischoff et al., 2019), feature extraction in software product lines (Bakar et al., 2015), feature location (Dit et al., 2013), and high-level mapping studies on crowdsourced feedback (Wang et al., 2019; Dąbrowski et al., 2022). However, these studies do not provide a unified view of how feature requests are managed throughout the software lifecycle.
@@ 3.2.2. Inclusion/Exclusion Criteria
To identify the most relevant studies on feature request analysis and processing, we drew upon prior work (Giray et al., 2023; Zakari et al., 2020; Croft et al., 2022; Niu et al., 2025b; Hu et al., [n. d.]) and established a set of inclusion (ICs) and exclusion criteria (ECs), summarized in Table 1. These criteria were applied to the titles, abstracts, and keywords of candidate papers to ensure relevance and quality.
@@ 3.3. Data Extraction and Synthesis
    • Data preprocessing techniques and vectorization methods are described.
@@
The goal of this survey is to aggregate evidence from primary studies and provide an overview of the state of the art on feature request analysis and processing. Following the systematic review guidelines (Keele et al., 2007), we synthesized the extracted data using a combination of descriptive (quantitative) synthesis and qualitative thematic synthesis, aligned with our four research questions. For RQ1 (Demographics), we performed descriptive synthesis on bibliographic and contextual attributes (e.g., publication year, venue type, and study type). We report the evolution of the literature over time and the distribution across venues using frequency counts. For RQ2 (Analysis and Processing Practices), we synthesized the extracted methodological information by grouping studies according to the main feature request tasks they target (e.g., classification, prioritization, summarization, management) and the approaches they employ. We summarize the prevalence of different practices (including preprocessing, vectorization/representation, and evaluation setups) using structured comparison tables, and we narratively discuss commonalities and differences across studies. For RQ3 (Public Datasets and Tools), we first summarized availability-related information (e.g., whether datasets/tools are publicly accessible, their sources) in a structured catalog. We then conducted a qualitative categorization of the datasets and tools by their intended purposes and research themes, and reported results for each category to facilitate reuse. For RQ4 (Challenges and Opportunities), we conducted an inductive thematic synthesis. During data extraction, any text explicitly describing a limitation, difficulty, research gap, or future work was copied verbatim into the extraction form. Two reviewers independently performed open coding on these excerpts to identify recurring concepts and iteratively refined a shared codebook through discussion. Conceptually similar codes were merged into higher-level themes, which were manually reviewed to ensure consistent granularity and clear boundaries. We report the resulting themes together with representative evidence from primary studies and, where appropriate, the number of studies mentioning each theme. When divergent or conflicting observations emerged, we explicitly described them rather than forcing a single conclusion.
@@ 4. Demographics
A total of 140 research papers were included in the final study set, covering publications up to August 2025. As shown in Figure 4, the number of primary studies has increased significantly over time, showing a clear upward trend from 2011 to 2017, with the peak occurring in 2017 and 2019 (16 studies). Activity remained relatively high through 2020, followed by a decline likely due to shifting research priorities. A notable resurgence occurred in 2024 (11 studies), coinciding with the rise of LLMs, which introduced new research opportunities. Overall, the data reveals cyclical research activity with notable peaks and dips over the years. Figure 5 presents the distribution of the selected papers across journals and conferences, highlighting the most frequently appearing publication venues. A total of 88 papers were published in conferences, 46 in journals, and 6 as preprints. Most studies were published in top-tier software engineering outlets, particularly in the field of RE. High-frequency venues include International Conference on Requirements Engineering (RE Conf., 15 papers), International Conference on Automated Software Engineering (ASE Conf., 8 papers), International Working Conference on Requirement Engineering: Foundation for Software Quality (REFSQ, 7papers), International Conference on Software Engineering (ICSE, 7 papers), Empirical Software Engineering (ESE, 7 papers), Journal of Systems and Software (JSS, 6 papers), Transactions on Software Engineering (TSE, 4 papers), International Conference on Software Maintenance and Evolution (ICSME, 4 papers), Requirements Engineering Journal (RE Jour., 4 papers), and International Conference on Evaluation and Assessment in Software Engineering (EASE, 4 papers). These results confirm that the topic has received strong attention from the software engineering community.
@@ 5.2. Quality Review
Quality review is a critical activity in RE, aimed at assessing the quality, correctness, and feasibility of incoming requirements. As noted by Wiegers (Wiegers, 2003), a comprehensive review should examine aspects such as the organization, completeness, correctness, quality attributes, and traceability of the requirements specification. Functioning as a “filter” within the requirements analysis process, feature request review helps eliminate low-quality requests—such as those that are ambiguous, redundant, or irrelevant—before they reach the development phase, thereby reducing errors and defects downstream (Wiegers, 2003).
@@ 5.4.1. Lexical Analysis-Based Approaches
Regarding priority ranking and competitive analysis, Licorish et al. (Licorish et al., 2017) focused on noun extraction from reviews to predict the processing priority of feature requests, while Dalpiaz et al. (Dalpiaz and Parente, 2019) utilized POS tagging and various POS combinations, such as <noun+adjective><noun+adjective> and <adjective+verb><adjective+verb>, to mine features from competitor app reviews. Shah et al. (Shah et al., 2018) proposed annotation guidelines emphasizing the labeling of only consecutive word phrases containing nouns, limited to a maximum of three words, while excluding references to the application itself.
@@ 5.7. Classification
✎ Classification Categories. Regarding classification categories, feature requests are typically organized according to requirement levels described in the requests, such as security, usability, ease of use, among others. RE commonly distinguishes between functional requirements—detailing software functionalities—and non-functional requirements, which specify quality attributes essential for fulfilling user and business needs. Standards like ISO 25010 categorize software quality into functional and non-functional characteristics, with the latter encompassing functional suitability, performance efficiency, compatibility, usability, reliability, security, maintainability, and portability (ISO/IEC, 2001). Various studies have classified feature requests accordingly. Li et al. (Li et al., 2018) categorized requests into security, reliability, performance, lifecycle, usability, ease of use, and system interface. Niu et al. (Niu et al., 2021) adopted the same dataset as Li et al. Scalabrino et al. (Scalabrino et al., 2019) categorized user comments into defect reports, functional requests, performance requests, security requests, energy consumption requests, availability requests, and other categories. Tao et al. (Tao et al., 2020) classified security-related comments into system, finance, spam, privacy, and others. Wang et al. (Wang et al., 2018) grouped requirements from user reviews into usability, reliability, portability, and performance. More recently, Bai et al. (Bai et al., 2024) used structural topic modeling to identify eight key user concerns from video conferencing app reviews, including data privacy, audio/video quality, customer service, security issues, meeting management, feature requests, and platform compatibility.
In conclusion, feature request classification is a crucial step in requirements analysis that facilitates effective organization, understanding, and management of diverse user requests. The field has evolved from early unsupervised clustering techniques to more advanced supervised learning and explainable AI approaches, leveraging both textual and auxiliary features for improved accuracy. Categorization schemes generally align with established RE frameworks, distinguishing functional and non-functional aspects and capturing detailed quality attributes. Continued advancements in classification methodologies and feature representation promise to further enhance automated analysis and prioritization of feature requests in software development.
@@ 5.9.1. Horizontal Traceability of Feature Requests
Horizontal traceability involves creating connections among software artifacts that exist at the same abstraction level or within a single development phase, particularly within RE (Heck and Zaidman, 2014). For feature requests, establishing horizontal traceability aids developers and stakeholders in gaining clear insights into project progress, efficiently managing the integration of new features, detecting requirement conflicts at an early stage, comprehending the rationale behind software requirements, handling interdependencies logically, and avoiding redundant or contradictory tasks, thereby reducing risks of software project failures.
@@ 5.11. Code Localization
One of the earliest and most influential studies in this field is by Marcus et al. (Marcus et al., 2004), who introduced an IR-based approach using LSI for concept and feature location. The method represents source code entities as textual documents and computes their semantic similarity with feature request descriptions in a latent space, enabling the identification of related code even without exact keyword matches. This seminal work laid the foundation for later IR-based and hybrid localization techniques.
@@ 6.2. Publicly Available Datasets
A significant impediment in many areas of RE research is the scarcity of publicly accessible datasets, often due to the proprietary nature of software requirements (Yu et al., 2024). However, the domain of feature requests presents a unique advantage: a substantial volume of data originates from open platforms, such as public repositories and app stores. This accessibility, often leveraged through web scraping techniques, has partially alleviated the data scarcity challenge that plagues other RE sub-fields. Consequently, various datasets have been curated and made publicly available by researchers. Table LABEL:tab:dataset_long catalogs these resources, detailing their origins and the specific research tasks they are typically used for (e.g., training classification models, evaluating duplicate detection).
@@
(Wang et al., 2023) https://github.com/KeyL99/VoteBot/ Identify feature requests from user review. 2023
@@
(Scalabrino et al., 2019) https://dibt.unimol.it/reports/clap/ Prioritization. 2017
@@
(Palomba et al., 2017) https://sites.google.com/site/changeadvisormobile/ Feature location. 2017
@@
(Hu et al., 2022) https://github.com/lcynju/review2code Feature location. 2022
@@ 7.1. Summarization
Before delving into the core findings, the study established essential context by defining software features, feature requests, and related concepts, underscoring their value and distinguishing them from traditional, formally elicited software requirements. We also positioned this review in relation to existing surveys, clarifying its unique contribution in providing an end-to-end RE perspective on the feature request lifecycle (Section 2).
The main body of our findings (presented from Section 4 onwards) begins by outlining the demographics of the research landscape. This revealed a consistent growth in publications dedicated to feature requests since 2010, with significant activity concentrated in reputable software engineering and RE venues, confirming the topic’s increasing importance within the community. We systematically cataloged the diverse origins of feature requests, confirming that while dedicated platforms like issue trackers are used, a vast amount originates from less structured channels such as app reviews, social media, and developer communications, with app reviews and issue reports being the most frequently studied sources (Section 5.1).
Addressing RQ3, the core analysis examined how these requests are analyzed and processed within a RE framework. We reviewed various automated methods designed to identify feature request descriptions from often noisy and informal user feedback, spanning unsupervised, supervised, and semi-supervised paradigms. Then we discussed techniques for software feature extraction, often relying on lexical approaches like POS tagging. It also included methods for structural analysis aimed at understanding content, such as user intent recognition, classification into diverse functional or non-functional categories, and automated summarization. Furthermore, the critical task of prioritization was examined, contrasting traditional RE approaches with methods adapted specifically for the high volume and nature of user feedback. Moreover, we examined approaches for quality review, primarily focusing on quality assessment related to duplication, redundancy, and ambiguity. We also explored traceability methods, encompassing both horizontal links between related requests and vertical links to artifacts such as source code and developers. Finally, various facets of change management were reviewed, including predicting request approval, recommending suitable developers, localizing relevant code sections, and suggesting APIs or refactorings.
@@ 7.2. Challenges and Opportunities
Comparative Analysis of Feature Requests and Bug Reports. Firstly, this review has focused exclusively on feature requests. However, user feedback often intertwines requests for new functionality with reports of existing problems (Dąbrowski et al., 2022). A crucial future direction is to integrate the study of bug reports alongside feature requests. Adopting a RE lens, comparative analyses are needed to explore the similarities and differences in effectively analyzing, prioritizing, and managing these two distinct yet interconnected types of feedback. Understanding where techniques converge and where they must diverge could lead to more holistic and efficient feedback management frameworks and tooling.
@@
Specification and Validation of Feature Requests. Furthermore, a significant and largely unaddressed gap identified in our study pertains to the specification and validation of feature requests. Currently, research heavily emphasizes identification and analysis, but largely overlooks the critical subsequent steps needed to bridge the gap between raw user input and actionable, verifiable requirements. Future research must urgently tackle the challenge of transforming collections of informal, often ambiguous, and fragmented user requests into more structured, consistent, and usable requirement representations. This could involve exploring techniques for semi-automatically generating preliminary user stories, use case scenarios, behavioral models, or even identifying constraints and non-functional aspects implicitly mentioned in feedback. Equally vital, and even less explored, is the validation of these derived specifications (Mallya et al., 2026). Methodologies and practical techniques are needed to ensure that the interpretations and specifications derived from user feedback accurately reflect the original user needs and intent before substantial development effort is invested. Addressing this specification and validation gap is fundamental to truly leveraging feature requests for building software that meets genuine user expectations. We plan to conduct in-depth investigations into these specific areas in our future work.
@@
Leveraging LLMs for Feature Request Intelligence. With the rapid advancement of LLMs, their application in software engineering tasks has gained increasing traction (Yu et al., 2024). However, our review finds that feature request research has yet to leverage the potential of LLMs fully. These models offer promising capabilities for a wide range of tasks, including classifying user intent, mapping feature requests to relevant source code locations, summarizing feedback, and recommending suitable developers or implementation strategies. Future work should explore how LLMs can be fine-tuned or prompted for domain-specific use in feature request processing, and investigate their effectiveness compared to traditional machine learning techniques. Moreover, new evaluation criteria and human-in-the-loop workflows may be necessary to ensure reliability, interpretability, and alignment with user expectations.
@@
Lack of Benchmarks for LLM-Driven Feature Request Tasks. Despite growing interest in applying LLMs to software maintenance and requirements tasks, there is a notable absence of standardized benchmarks specifically targeting LLM-based feature request processing (Hu et al., [n. d.]). This limits reproducibility, fair comparison across methods, and progress tracking over time. To address this, the research community should work toward building and releasing well-curated, multi-task benchmark datasets that reflect real-world feature request scenarios. These benchmarks should include annotations for classification, prioritization, code mapping, and developer assignment, and be accompanied by baseline LLM implementations to support further development and evaluation.
@@
Multilingual and Cross-Cultural Feature Request Processing. Current research on feature requests is heavily skewed toward English-language feedback (Bhuiyan et al., 2026), predominantly from Western platforms and user bases. However, in today’s global software market, products often receive feedback in multiple languages and cultural contexts. This introduces new challenges in understanding linguistic nuances, implicit user expectations, and regional preferences. Future work should explore multilingual models and culturally adaptive techniques to support accurate identification and interpretation of feature requests across diverse user groups. This direction also calls for the development of multilingual corpora and evaluation tools.
@@
Bridging User and Developer Perspectives. Feature requests often originate from non-technical users but are consumed and implemented by developers, creating a semantic and expectation gap. Bridging this gap remains an underexplored area. Future work could investigate how to translate user-facing language into developer-understandable representations, integrate user sentiment and urgency into development planning, and develop tools that facilitate two-way communication between stakeholders. Enhancing alignment between user needs and developer actions could significantly improve the effectiveness and user satisfaction of feature implementation (Bano et al., 2017).
Autonomous Agents for End-to-End Feature Request Processing. Another emerging and promising direction for future research is the development of autonomous agents capable of handling feature requests throughout the entire lifecycle, from initial user interaction and intent interpretation to analysis, code localization, and implementation. With advancements in multi-agent collaboration frameworks and AI planning (Zhang et al., 2025; Tran et al., 2025), it is increasingly feasible to envision intelligent agents that can process user feedback, generate actionable development tasks, and even coordinate with other agents for implementation and testing. Such agent-based systems could significantly reduce the human effort required for managing feature requests, especially in large-scale and user-driven software systems. Exploring the design, capabilities, and limitations of such autonomous feature request agents offers a valuable avenue for advancing intelligent software engineering.